Can someone guide me on how to do this? I’m not quite understanding what happens and where the electrons go when the spheres touch and why. Thank you !

5. CESİ] Consider three identical metal spheres, A, B, and C Sphere A carries a charge of +5q. Sphere B carries a charge of -q. Sphere C carries no net charge. Spheres A and B are touched together and then separated. Sphere C is then touched to sphere A and separated from it. Last, sphere C is touched to sphere B and separated from it. (a) How much charge ends up on sphere C? What is the total charge on the three spheres (b) before they are allowed to touch each other and (c) after they haye touched?
Answer #1

Note: whenever the spheres will touch, the charge on them will be equally divided. But total will be equal to total initial charge.

A and B are touched.
Total charge = 5 + (-1) = 4q. Both will get equal charge.
Now A and B both have 2q

Now C is touched to A
Total charge = 0 + 2 = 2q.
Now Both have 1q each

Now C is touched to B
Total charge = 1 + 2 = 3q
Now both have 1.5q

Finally

a)

A = 1q; B = 1.5q ; C = 1.5q

b)

Total initial = 5 + -1 +0= 4q

c)

Total final = 1 + 1.5 + 1.5 = 4q

*No charge is lost. This is called law of conservation of charge.